OpenStudy (anonymous):

mary reduced the size of a painting to a width of 3.3 in. what is the new height if it was originally 32.5 in tall and 42.9 in wide?

OpenStudy (zpupster):

this is a ration problem height/width= new ht /new wdth height=32.5 width=42.9 new ht =x new wdth=3.3 i use height instead of tall and watch your units in these kind of problems so plug in your values 32.5/42.9=x/3.3 cross multiply 3.3(32.5)=42.9x divide both sides by 42.9 isolating x 3.3(32.5)/42.9 =x x=???